# Language Bottleneck Models: A Framework for Interpretable Knowledge Tracing and Beyond

### v1

@@ B.5 Example bottlenecks on the Eedi dataset
Figure B2 shows example bottlenecks produced by a gpt-4o-based LBM for a student in the Eedi dataset. The input data is composed of 30 questions from various constructs, and the decoder predicts 4 test questions. As shown Figure 4 in the main paper, a shorter bottleneck of 128 tokens constrains the expressive power of the model, and in this example the resulting predictions fail on two of the test questions. A longer bottleneck of 256 or 512 tokens allows for more nuance and details in the bottleneck, leading to the decoder correctly predicting all four test questions for this student.
@@
Figure B2: Example bottlenecks produced by an LBM with gpt-4o as both encoder and decoder for a student in the Eedi dataset, with varying bottleneck lengths. The input data provided to the encoder is composed of 30 questions across various constructs, and the decoder predicts 4 test questions using the bottleneck.
Appendix C Extended Discussion
C.1 Extended Discussion
Knowledge Tracing as an Inverse Problem
The reframing of Knowledge Tracing as an inverse problem represents a fundamental shift in how we conceptualize student modeling. Traditional KT methods adopt a forward modeling approach, where the goal is to learn a mapping from observed interactions to future performance predictions. While effective for prediction, this treats the student’s knowledge state as an opaque latent variable that serves primarily as an intermediate computational step. In contrast, our inverse problem formulation explicitly seeks to recover the underlying knowledge state that generates observed behaviors, making interpretability the primary objective rather than an afterthought. This aligns with educational assessment principles, where the main objective is to understand what a student knows, rather than merely predicting future behavior.
Parallel with Kolmogorov Complexity
Our inverse problem formulation of Knowledge Tracing draws a natural parallel with algorithmic information theory. The Kolmogorov complexity $K(x)$ of a string $x$ is defined as the length of the shortest program that outputs $x$ when run on a universal Turing machine. Traditional KT methods focus on learning statistical regularities between questions and responses, without looking for parsimonious explanations for observed behavior. In contrast, LBMs explicitly search for minimal natural language descriptions that can both reconstruct past interactions and predict future responses. The knowledge state summary $\mathcal{S}$ is akin to the minimal program, and the decoder LLM functions is akin to the universal interpreter that unpacks $\mathcal{S}$ to generate the observed question-answer patterns. This connection suggests that effective knowledge state summaries should capture the algorithmic essence of student behavior—the underlying "program" of knowledge and misconceptions that generates observable responses—rather than merely fitting surface-level patterns. While true Kolmogorov complexity is uncomputable, LBMs approximate this ideal through the natural language bottleneck constraint, encouraging summaries that balance compression with predictive power.
Computational Architecture and Design Choices
C.2 Extended Limitations
Context Length Constraints
Computational Cost and Resource Requirements
Dependence on Textual Question Content
Many existing KT datasets provide only question identifiers rather than full question text. This limits the direct applicability of LBMs, as they require the questions content in order to generate meaningful summaries about student knowledge. While question content is increasingly available in modern educational platforms, this dependency creates a barrier to applying LBMs to historical datasets or systems that rely primarily on item response theory frameworks.
Quasi-Knowledge State Assumption
The quasi-static knowledge state assumption underlying our inverse problem formulation, while reasonable for short diagnostic sessions, does not cover longer time horizons. Real learning involves continuous knowledge acquisition, forgetting, and misconception evolution that our current framework cannot capture. This limitation restricts current LBMs to diagnostic rather than contexts of assessment throughout the learning process. The assumption also fails to account for contextual factors (fatigue, motivation, external stressors) that can significantly impact student performance within even short sessions.
C.3 Extended Future Work
Iterative Encoding Strategies
Active Learning and Question Selection
Dynamic Knowledge State Modeling
Extending LBMs to handle evolving knowledge states represents a significant next step necessary to account for progressive learning and forgetting effects. A natural relaxation of the static knowledge state assumption involves "piecewise-constant" knowledge states that can evolve between question sessions but remain static within sessions. This extension poses exciting challenges in developing training objectives that balance within-session consistency with between-session learning dynamics, potentially requiring new approaches to temporal modeling in natural language representations.
Multimodal Input Integration
Pedagogical Alignment
Recent work like LearnLM [28] demonstrates the potential for making LLMs more pedagogically aligned. Incorporating similar pedagogical principles into LBM components could help encoders better interpret educational interactions and generate summaries that align with expert teaching practices. This might involve training on educator-annotated examples, incorporating educational taxonomies into summary structure, or using reward functions that emphasize pedagogically relevant aspects of student knowledge states. Such alignment could bridge the gap between computational convenience and educational validity.
Appendix D Experimental Details
D.1 Models
LLM models
@@ D.1 Models
  • GPT-4o and GPT-4o-mini[2];
  • Qwen2.5-3B and Qwen2.5-7B[29];
  • Gemma3-12B and Gemma3-27B[27].
For open-source models we extract the activation logits of the “Yes” and “No” tokens and return the higher value. For closed-source models we prompt for a “Yes” or “No” answer and parse the text output.
@@
We evaluate the following four Knolwedge Tracing models:
  • DKT [20]
  • DKVMN [34]
  • AKT [6]
  • SAKT [19]
For each one of these models, we use the pyKT implementation [15] with default hyperparameters.
D.2 Datasets
D.2.1 Synthetic
@@
  • fails whenever an operand $>10$;
@@
Figure D1 shows the histogram of the number of misconceptions per student across the $2000$ students of the dataset.
Figure D1: Distribution of the number of misconceptions per student in the Synthetic dataset.
D.2.2 Eedi
The Eedi dataset is analogous to the one publicly shared via the NeurIPS 2020 Education Challenge [30] organised by Eedi, but additionally includes the text of each question.
Although many trajectories in the Eedi dataset span several days or months, we retain only single-session sequences to satisfy the quasi-static-knowledge assumption:
  • minimum response time: $5\,\mathrm{s}$;
@@ D.2.2 Eedi
This filtering yields a total of $623$ individual trajectories with 40+ questions.
D.3 Results
D.3.1 Table 1
Synthetic dataset
  • $|\mathcal{X}|=50$
  • $|\mathcal{Y}|=4$
  • Test students: $200$
  • Bottleneck size: $128$ tokens
  • Chain-of-thought prompting: No
Eedi dataset
  • $|\mathcal{X}|=30$
  • $|\mathcal{Y}|=4$
  • Test students: $100$
  • Bottleneck size: $512$ tokens
  • Chain-of-thought prompting: No
D.3.2 Figure 4
Same set-up as Table 1; only the bottleneck size varies (see $x$-axis).
D.3.3 Figure 5
@@
  • Chain-of-thought prompting: Yes
  • Encoder & decoder: identical (varies by row in the table)
Knowledge-tracing baselines DKT [20], DKVMN [34], AKT [6], and SAKT [19] are evaluated using the pyKT implementation [15], modified to compute accuracy on $|\mathcal{Y}|=4$ questions, in order to be comparable with our LBM models. For each method we keep the pyKT default hyperparameters.
D.3.4 Figure 6
@@
  • Learning $\text{Learning rate}=1\times 10^{-4}$
@@
  • Optimiser: Adam (adamw_torch, default settings)
@@
D.3.5 Encoder–Decoder Variants
Identical to the Table 1 set-up, except for the choice of encoder and decoder LLMs.
D.3.6 Difficulty Variations
Synthetic students are stratified by the number of misconceptions held: 00, $1$, $2$, or $3+$.
D.3.7 Steering Experiments
@@
  • Learning $\text{Learning rate}=1\times 10^{-4}$
@@
  • Optimiser: Adam (adamw_torch, default settings)
@@ 2 Here is some information about a student:
5 {y_prompt}
@@ 8 Input:
Listing 3: Base Prompt for the Encoder when steering for mentioning "misconceptions" (Section B.4.1).
D.5 Code and hardware

### v2

@@ A.6 Example bottlenecks on the Eedi dataset
Figure A4 shows example bottlenecks produced by a gpt-4o-based LBM for a student in the Eedi dataset. The input data is composed of 30 questions from various constructs, and the decoder predicts 4 test questions. As shown Figure 4 in the main paper, a shorter bottleneck of 128 tokens constrains the expressive power of the model, and in this example the resulting predictions fail on two of the test questions. A longer bottleneck of 256 or 512 tokens allows for more nuance and details in the bottleneck, leading to the decoder correctly predicting all four test questions for this student.
@@
Figure A4: Example bottlenecks produced by an LBM with gpt-4o as both encoder and decoder for a student in the Eedi dataset, with varying bottleneck lengths. The input data provided to the encoder is composed of 30 questions across various constructs, and the decoder predicts 4 test questions using the bottleneck.
Appendix B Experimental Details
B.1 Language Bottleneck Models
LLM backbones
@@ B.1 Language Bottleneck Models
  • GPT-5, GPT-4o and GPT-4o-mini(Achiam et al., 2023);
  • Qwen2.5-3B and Qwen2.5-7B(Team, 2024);
  • Gemma3-12B and Gemma3-27B(Team, 2025).
For open-source models we extract the activation logits of the “Yes” and “No” tokens and return the higher value. The logits for the “Yes” and “No” tokens are used to compute the AUC. For closed-source models we prompt for a “Yes” or “No” answer and parse the text output. We run GPT models via the OpenAI API, with the following snapshots: GPT-5: 2025-08-07; GPT-4o: 2024-08-06; GPT-4o-mini: 2024-07-18.
Reward function
For training the encoder in Section 5, we set the reward function to the decoder accuracy across $|\mathcal{Y}|=20$ questions:
$R(\tilde{\mathcal{S}};g)=\texttt{Acc}\big(\tilde{\mathcal{Y}},\mathcal{Y}\big)$
The RL steering experiment section A.4.2 only rewards the presence of the word "misconception" in the bottleneck:
$R(\tilde{\mathcal{S}};g)=\Omega(\tilde{\mathcal{S}})=\mathbbm{1}[\text{'misconception'}\in\tilde{\mathcal{S}}$]
B.2 Training the decoder
To evaluate the performance of a trained decoder (as shown in Figure 6 and Table A.2), we fine-tune a gemma-3-12b-it decoder models based on a frozen gpt-5 encoder on both EEDI and XES3G5M.
Data Construction
We train on 800 (resp. 400) students from the training split of the XES3G5M Filtered (resp. EEDI Filtered) dataset, which amounts to about 80% of the data, leaving 200 test students. For each student interaction history $H_{t}$, we first generate a bottleneck summary $\tilde{\mathcal{S}}=f_{\theta}(H_{t})$ using the frozen gpt-5 encoder. We then pair this summary with the ground-truth binary correctness labels for the $|\mathcal{Y}|=4$ target questions associated with that student, yielding a training dataset of $N=3,200$ (resp. $N=1,600$) training examples.
Training Procedure
For each bottleneck $\tilde{\mathcal{S}}$ and question $q$, the model is prompted to predict the correctness of the student via the decoder prompt shown in Listing LABEL:lst:base_prompt_dec. We then train with SFT on the ground truth answers using LoRA (Hu et al., 2022) on a single NVIDIA RTX 6000 48GB GPU. We use a train/validation split of 90/10 and train with early stopping on validation accuracy computed after each epoch.
The training configuration:
  • Model: google/gemma-3-12b-it
  • LoRA Configuration: Rank $r=16$, Alpha $\alpha=16$. Target modules: [q_proj, k_proj, v_proj, o_proj].
  • Optimization: AdamW optimizer with a learning rate of $1\times 10^{-4}$.
  • Batch Size: Effective batch size of 8 (per-device batch size of 2 with 4 gradient accumulation steps).
  • Duration: The model was trained for up to 5 epochs with early stopping after the first decreasing validation accuracy.
B.3 KT/CD Models
Cognitive Diagnosis models
We evaluate the following 5 Cognitive Diagnosis models:
  • IRT (Rasch, 1993)
  • MIRT (Reckase, 2006)
  • DINA (Junker and Sijtsma, 2001)
  • KaNCD (Wang et al., 2022)
  • NeuralCDM (Wang et al., 2022)
We use the implementation from the EduCDM library (bigdata-ustc, 2021). To make sure there is enough data per student to train on, we filter out students with <10<10 interactions in each dataset.
CD baselines CD models were implemented using the EduCDM library (bigdata-ustc, 2021). Unlike KT and LBMs, CD models learn static, student-specific parameters (e.g., proficiency embeddings) which requires every student to be present during training. Consequently, they cannot be evaluated on previously unseen students in a zero-shot manner. To ensure a fair comparison, we evaluate CD models using a held-out set of $|\mathcal{Y}|=4$ interactions for each student. In this setup, the model “reads” the student’s history by optimizing student-specific parameters via gradient descent. We consider this functionally equivalent to how KT and LBMs process the same history “in-context” (via RNN states or attention) to adapt to a new student. Therefore, despite the protocol difference, all models use the same historical information $\mathcal{X}$ to predict $\mathcal{Y}$.
@@ B.3 KT/CD Models
We evaluate the following 9 Knowledge Tracing models:
  • DKT (Piech et al., 2015)
  • DKVMN (Zhang et al., 2017)
  • SAKT (Pandey and Karypis, 2019)
  • GKT (Nakagawa et al., 2019)
  • Deep IRT (Yeung, 2019)
  • AKT (Ghosh et al., 2020)
  • SAINT (Choi et al., 2020)
  • SimpleKT (Liu et al., 2023a)
  • QIKT (Chen et al., 2023)
KT baselines KT models are evaluated using the pyKT library (Liu et al., 2022), modified to compute accuracy on the same $|\mathcal{Y}|=4$ target window used for LBMs. We use the default hyperparameters from pyKT for all models. Test accuracy is computed on a set of $200$ unseen test students.
B.4 Datasets
B.4.1 Synthetic
@@
  • fails whenever an operand >10>10;
@@
Figure B1 shows the histogram of the number of misconceptions per student across the $2000$ students of the dataset.
Figure B1: Distribution of the number of misconceptions per student in the Synthetic dataset.
B.4.2 Eedi
The Eedi dataset is analogous to the one publicly shared via the NeurIPS 2020 Education Challenge (Wang et al., 2020) organised by Eedi, but additionally includes the text of each question. While the exact version of the dataset used in this work is not available publicly, a very similar version including question texts is available via the "Eedi - Mining Misconceptions in Mathematics" Kaggle Competition (King et al., 2024).
Preprocessing
Although many trajectories in the Eedi dataset span several days or months, we retain only single-session sequences to satisfy the constant knowledge state assumption - a common assumption in Cognitive Diagnosis (Wang et al., 2024). We compute sessions of interactions by grouping answers that are within the follow criteria:
  • minimum response time: $5\,\mathrm{s}$, to avoid random answers;
@@ B.4.2 Eedi
This filtering yields the "EEDI Filtered" dataset with a total of $623$ individual trajectories with 40+ questions.
B.4.3 XES3G5M
The XES3G5M dataset contains student interaction data from a large-scale online mathematics learning platform (Liu et al., 2023b). The original dataset has question and construct text in Chinese. We use English translation from Ozyurt et al. (2024) to also run on an English version of the dataset.
Preprocessing
The XES3G5M preprocessing follows a similar approach to Eedi, with adjusted parameters to accommodate the characteristics of this dataset. We compute sessions of interactions by grouping answers that meet the following criteria:
  • minimum response time: $5\mathrm{s}$, to avoid random answers;
  • maximum gap between answers: $10\mathrm{min}$;
  • minimum trajectory length: $34$ questions.
The increased maximum gap between answers (10 minutes vs. 3 minutes for Eedi) is to ensure that a sufficient number of sessions are available for training. More stringent filtering make it easier to satisfy the constant knowledge state assumption, but might not produce enough data points for effective training of KT and CD models. This filtering yields the "XES3G5M Filtered" dataset with a total of $1,245$ student-sessions trajectories with 34+ questions, across $996$ individual students.
B.5 Experiment Parameters
B.5.1 Table A.2, Figure 6A, Table 2, Figure 7
Table B1 summarize experimental settings used for Table A.2, Figure 6, Table 2 and Figure 7. Runs are aggregated across N=3 for LBMs/Direct LLMs and N=10 for KT/CD models. For Figure 7 the bottleneck size varies along the plot’s x axes.
Table B1: Experimental settings for the different datasets.
Synthetic Eedi (Filtered) XES3G5M (Filtered)
$|\mathcal{X}|$ 50 30 30
$|\mathcal{Y}|$ 4 4 4
# Train students 400 800
# Test students 200 100 200
Bottleneck size 128 tokens 512 tokens 512 tokens
CoT prompting No No No
Statistical Test for Table 2
Table 2 shows in bold the models and datasets for which the LBM variant has an accuracy of no more than 2% below the Direct baseline. We use Welch’s t-test with the null hypothesis: $\texttt{accuracy}_{LBM}>=\texttt{accuracy}_{direct}-0.02$‘, and report in bold runs where $p<0.05$.
B.5.2 Figure 6B
@@
  • Chain-of-thought prompting: Yes; reasoning_effort=minimal for GPT-5
  • Encoder & decoder: same backbone LLM (varies by row in the table)
B.5.3 Training the encoder (Figure 5)
@@
  • $\text{Learning rate}=1\times 10^{-4}$
@@
  • Optimizer: Adam (adamw_torch, default settings)
@@
B.5.4 Encoder–Decoder Variants
Identical to the Table A.2 set-up, except for the choice of encoder and decoder LLMs.
B.5.5 Steering Experiments
@@
  • $\text{Learning rate}=1\times 10^{-4}$
@@
  • Optimizer: Adam (adamw_torch, default settings)
@@ 6 Input:
12
13# with CoT
14cot_instruction: |
15 Think step by step and lay out your reasoning before you write the final summary. Then, enclose the final summary in <info>...</info>.
16# without CoT
17cot_instruction: |
18 Enclose your entire summary in <info>...</info> and do not include anything else.
19
20input_text: |
21 The student answered the following questions:
22 {question_1_text}
23 ...
24 {question_n_text}
25
26# with construct information
27question_i_text: |
28 Question {question_txt, question_ID}, with construct{construct_txt, construct_ID}: answered {correctness}.
29
30# without construct information
31question_i_text: |
32 Question {question_txt, question_ID}: answered {correctness}.
@@ 2 Here is some information about a student:
5 Predict whether the student will answer {new_question} correctly or not.
6 Answer with "Yes" or "No" and nothing else.
@@ 8 Input:
Listing 3: Base Prompt for the Encoder when steering for mentioning "misconceptions" (Section A.4.2).
B.7 Code and hardware